FlowerMate is a manufacturer of large flower pots for urban settings. The company has these​ standards:

Direct materials (resin) - 15 pounds per pot at a cost of $6.00

Direct labor - 2.0 hours at a cost of $16.00 per hour

Standard variable manufacturing overhead rate - $7.00 per direct labor hour

Budgeted fixed manufacturing overhead - $20,600

Standard fixed MOH rate - $6.00 per direct labor hour (DLH)

FlowerMateFlowerMate

allocated fixed manufacturing overhead to production based on standard direct labor hours. Last​ month, the company reported the following actual results for the production of 1,800

flower​ pots:

Direct materials - Purchased 28,480 pounds at a cost of $6.50 per pound; used 28,080 pounds to produce 1,800 pots.

Direct labor -Worked 2.5 hours per flower pot (4,500 otal DLH) at a cost of $14.00 per hour

Actual variable manufacturing overhead- $7.40 per direct labor hour for total actual variable manufacturing overhead of $33,300

Actual fixed manufacturing overhead - $19,900

Standard fixed manufacturing overhead allocated based on actual production- $21,600

Requirements

1. Compute the direct material price variance and the direct material quantity variance.

2. Who is generally responsible for each​ variance?

3. Interpret the variances.
Answer #1

1. Direct material price variance = $11240 (unfavorable or adverse)

Direct material quantity variance = $6480 (unfavorable or adverse)

2. The Purchasing or Procurement team is responsible for Direct material price variance since its is their responsibility to buy products at a given price.

The production team primarily the production managers are responsible for Direct material quantity variance since its is their responsibility to ensure that only the given quantity of material is used during the production process.

3. Direct material price variance shows the variance arising on account of difference between standard and actual price of goods procured. In the given question, the company end up spending $0.5 additional per pound over and above the standard price resulting in unfavorable price variance.

Direct material quantity variance shows the variance arising on account of difference between standard and actual quantity of direct material used in production. In the given question, the company ended up spending 1080 pounds additional to produce the desired quantity. This resulted in in unfavorable quantity variance.

Notes

Direct Material Price Variance = (Standard Price - Actual Price)*Actual Quantity Purchased

=($6 - $6.50) * 28480 = $11240

Direct Material Quantity Variance = (Standard Quantity - Actual Quantity)*Standard Price

= (15*1800 - 28080) * 6 = $6480

Please comment in case of any issue and I will be happy to help.